•The half-life period of yttrium is 20 min. If there are 1000 g of yttrium present, how much will be left after 30 min?
1
Expert's answer
2020-06-22T11:09:16-0400
"m=M\cdot 2^{-\frac{t}{T}}\\m=1000\cdot 2^{-\frac{30}{20}}=354\ g"